# Improved Regret Bounds for Gaussian Process Upper Confidence Bound in Bayesian Optimization

### v2

@@ A.2 Proof of Lemma 5
By aggregating the conditions (82), (92), and (98), we set $\overline{T}$ as the smallest natural number such that the following inequalities hold:
$\overline{T}$ $\geq\max\left\{\left(\frac{e^{2}c_{d}\widetilde{C}_{\mathrm{SE}}}{2\ell^{2}c_{\mathrm{quad}}}\right)^{2},\frac{e^{2}c_{d}\widetilde{C}_{\mathrm{SE}}^{2}}{2\ell^{2}c_{\mathrm{lin}}^{2}}\right\}\beta_{T}\ln^{d+1}T,~{}~{}\mathrm{and}$ (101)
$\overline{T}$ $\geq\max\left\{\left(\frac{ec_{d}}{2c_{\mathrm{quad}}\ell^{2}}\right)^{4},~{}\left(\frac{ec_{d}}{2c_{\mathrm{lin}}^{2}\ell^{2}}\right)^{2}\right\}\widetilde{C}_{\mathrm{SE}}^{4}\beta_{T}^{2}\ln^{2(d+1)}T.$ (102)
Then, from Eqs. (85), (87), (94), and (100), we have
$\sqrt{\max\left\{\gamma_{(T/2^{i-1})}(\mathcal{B}_{2}(\eta_{i}c_{\mathrm{lin}}^{-1};\bm{x}^{\ast})),\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)\right\}}=O(\sqrt{\ln T}).$ (103)
@@ A.3 Proof of Lemma 6
From the upper bound of the discretization error in event $\mathcal{A}$, we have $\forall t\geq\sqrt{2/\varepsilon},\forall\bm{x}\in\mathcal{X},|f(\bm{x})-f([\bm{x}]_{t})|\leq\varepsilon/2$. Here, we set $\underline{\mathcal{T}}(\varepsilon)=\{t\in\mathbb{N}_{+}\mid t\geq\sqrt{2/\varepsilon}\}$. By relying on the standard argument of MIG [48], we observe the following inequality for any realizations and $\varepsilon>0$:
@@
where the second inequality follows from the definition of $\mathcal{T}(\varepsilon)$, and the last inequality follows from $\widetilde{t}\in\underline{\mathcal{T}}(\varepsilon)$ and event $\mathcal{A}$. Therefore, under $\mathcal{A}$, the inequality $-\frac{\varepsilon}{2}+2\sqrt{\frac{C\beta_{\widetilde{t}}\gamma_{|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|}(\mathcal{X})}{|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|}}\geq 0$ must hold; otherwise, $\mu(\bm{x}_{\widetilde{t}};\mathbf{X}_{\widetilde{t}-1};\bm{y}_{\widetilde{t}-1})+\beta_{\widetilde{t}}^{1/2}\sigma(\bm{x}_{\widetilde{t}};\mathbf{X}_{\widetilde{t}-1})<\mu([\bm{x}^{\ast}]_{\widetilde{t}};\mathbf{X}_{\widetilde{t}-1};\bm{y}_{\widetilde{t}-1})+\beta_{\widetilde{t}}^{1/2}\sigma([\bm{x}^{\ast}]_{\widetilde{t}};\mathbf{X}_{\widetilde{t}-1})$, which contradicts $\bm{x}_{\widetilde{t}}\in\operatornamewithlimits{argmax}_{\bm{x}\in\mathcal{X}}\mu(\bm{x};\mathbf{X}_{\widetilde{t}-1};\bm{y}_{\widetilde{t}-1})+\beta_{\widetilde{t}}^{1/2}\sigma(\bm{x};\mathbf{X}_{\widetilde{t}-1})$. This further implies
@@ Appendix B Information Gain Upper Bound
Our analysis requires the upper bound of MIG with explicit dependence on the radius of the input domain. Several existing works [4, 25, 26] established such a result by extending the proof in [48]. However, the proof strategy in [48] result in $\widetilde{O}(T^{\frac{d(d+1)}{2\nu+d(d+1)}})$ upper bound of MIG in Matérn kernel, which is strictly worse than the best achievable $\widetilde{O}(T^{\frac{d}{2\nu+d}})$ upper bound. Vakili et al. [54] shows $\widetilde{O}(T^{\frac{d}{2\nu+d}})$ upper bound of MIG with $\nu>1/2$ under the uniform boundness assumption of the eigenfunctions. Furthermore, the following work [51] shows $\gamma_{T}(\{\bm{x}\in\mathbb{R}^{d}\mid\|\bm{x}\|_{2}\leq\eta\})=\widetilde{O}(\eta^{\frac{2\nu d}{2\nu+d}}T^{\frac{d}{2\nu+d}})$ for any radius $\eta>0$ if there exists eigenfunctions uniformly bounded without depending on $\eta>0$. Some of the related results supports the uniform boundness assumption under $d=1$ [25, 58], or under the approximated version of the original Matérn kernel [39, 47]; however, to our knowledge, we are not aware of any literature that rigorously support uniform boundness assumption under the general compact input domain with $d\geq 2$ and $\nu>1/2$. See Chapter 4.4 in [25] for the detailed discussion. Therefore, this section’s goal is twofold: (i) prove $\widetilde{O}(T^{\frac{d}{2\nu+d}})$ upper bound as claimed in [54] without relying on the uniform boundness assumption, and (ii) clarify the explicit dependence on the input radius in the upper bound proved in (i).
@@
While the above results ignore the explicit dependence on $d$ and $\nu$, all the other parameters are explicitly stated in our upper bound of MIG. We would like to emphasize that we do not rely on the uniform boundness assumption of the eigenfunctions to prove the above results. In the above results for $k=k_{\mathrm{SE}}$, we obtain the same $O(\ln^{d+1}T)$ upper bound as that in [54] except for the constant factor. For $k=k_{\mathrm{Mat\acute{e}rn}}$, we also obtain the same $\widetilde{O}(T^{\frac{d}{2\nu+d}})$ upper bound as that in [54], while the logarithmic dependence get worse from $O\left(\ln^{d/(2\nu+d)}T\right)$ to $O\left(\ln^{(4\nu+d)/(2\nu+d)}T\right)$. Furthermore, the above result reveals the explicit dependence of the radius $\eta$ of the input domain. Regarding the case $k=k_{\mathrm{Mat\acute{e}rn}}$, our result suggests $\widetilde{O}(\eta^{\frac{2\nu d}{2\nu+d}}T^{\frac{d}{2\nu+d}})$ upper bound of MIG, which is consistent with that in [51] with uniform boundness assumption.
@@
The basic proof strategy follows that in [54], which leverages the Mercer decomposition of the kernel. To bypass the uniform boundness assumption in the proof of [54], we must resort to other specific properties of the eigenfunction. However, except for some exceptional cases, the eigenfunction of the kernel on the general compact domain is difficult to specify in an analytical form and complex to analyze. To avoid this issue, instead of studying the original definition of the MIG on $\mathbb{R}^{d}$, we consider reducing the original MIG on $\mathcal{X}\coloneqq\{\bm{x}\in\mathbb{R}^{d}\mid\|\bm{x}\|_{2}\leq 1\}$ to that on a hypersphere $\mathbb{S}^{d}\coloneqq\{\bm{x}\in\mathbb{R}^{d+1}\mid\|\bm{x}\|_{2}=1\}$ defined in $\mathbb{R}^{d+1}$. The eigensystems on $\mathbb{S}^{d}$ are one of the exceptional cases, whose eigenfunctions are specified as a special function on $\mathbb{S}^{d}$, called spherical harmonics [1, 15]. Indeed, by using the addition theorem of the spherical harmonics (Theorem 14), the existing works [23, 29, 53] already demonstrated that the upper bound of MIG on $\mathbb{S}^{d}$ is proved as with [54] without the uniform boundness assumption. We use their technique to show the upper bound of MIG under SE and Matérn kernels on $\mathbb{R}^{d}$, while the original motivation of these existing works is to study the MIG under the neural tangent kernel on a hypersphere. The remaining parts of this section are constructed as follows:
@@
  • In Section B.3, we provide the general upper bound of the information gain on $\mathbb{S}^{d}$ (Lemma 15), represented by the kernel function’s eigenvalues. This subsection’s result has no intrinsic change from those in [29, 53]; however, we provide details for completeness.
@@ B.1 Reduction of the MIG on $\mathbb{R}^{d}$ to $\mathbb{S}^{d}$
where $[\widetilde{\mathbf{K}}]_{i,j}=k(\widetilde{\bm{x}}_{i},\widetilde{\bm{x}}_{j})/k(\bm{x}_{i},\bm{x}_{j})$. Here, $A\odot B$ denotes the Hadamard product of the matrices $A$ and $B$. Then, Oppenheim inequality (e.g., Theorem 7.27 in [59]) implies $\det(A\odot B)\geq\det(B)\prod_{i}A_{ii}$ for any positive semi-definite matrices $A$ and $B$. Therefore, if $\widetilde{\mathbf{K}}$ is a positive semi-definite matrix, Eq. (136) immediately implies
@@
Similarly to the proof for $k=k_{\mathrm{SE}}$, we consider the application of Oppenheim inequality; however, the positive semi-definiteness of element-wise quotient matrix $\widetilde{\mathbf{K}}$ is unknown for $k=k_{\mathrm{Mat\acute{e}rn}}$. To avoid this problem, we leverage the following representation of $k_{\mathrm{Mat\acute{e}rn}}$, which is given as the form of the lengthscale mixture of the SE kernel [50]:
@@
$k_{1}(\bm{x},\widetilde{\bm{x}})$ $=\frac{1}{\Gamma(\nu)}\int_{0}^{\eta_{1}}z^{\nu-1}e^{-z}\exp\left(-\frac{\|\bm{x}-\widetilde{\bm{x}}\|_{2}^{2}}{2\ell^{2}z\nu^{-1}}\right)\mathrm{d}z,$ (142)
$k_{2}(\bm{x},\widetilde{\bm{x}})$ $=\frac{1}{\Gamma(\nu)}\int_{\eta_{1}}^{\eta_{2}}z^{\nu-1}e^{-z}\exp\left(-\frac{\|\bm{x}-\widetilde{\bm{x}}\|_{2}^{2}}{2\ell^{2}z\nu^{-1}}\right)\mathrm{d}z,$ (143)
$k_{3}(\bm{x},\widetilde{\bm{x}})$ $=\frac{1}{\Gamma(\nu)}\int_{\eta_{2}}^{\infty}z^{\nu-1}e^{-z}\exp\left(-\frac{\|\bm{x}-\widetilde{\bm{x}}\|_{2}^{2}}{2\ell^{2}z\nu^{-1}}\right)\mathrm{d}z$ (144)
@@
$|k_{1}(\bm{x},\widetilde{\bm{x}})|$ $=\frac{1}{\Gamma(\nu)}\int_{0}^{\eta_{1}}z^{\nu-1}e^{-z}\exp\left(-\frac{\|\bm{x}-\widetilde{\bm{x}}\|_{2}^{2}}{2\ell^{2}z\nu^{-1}}\right)\mathrm{d}z$ (147)
@@
$|k_{3}(\bm{x},\widetilde{\bm{x}})|$ $=\frac{1}{\Gamma(\nu)}\int_{\eta_{2}}^{\infty}z^{\nu-1}e^{-z}\exp\left(-\frac{\|\bm{x}-\widetilde{\bm{x}}\|_{2}^{2}}{2\ell^{2}z\nu^{-1}}\right)\mathrm{d}z$ (156)
@@
where $\mathbf{K}_{2}^{(q)}(\mathbf{X}_{T},\mathbf{X}_{T})$ is the kernel matrix of $k_{2}^{(q)}$. Next, for any $q$, we define new kernel function $\widetilde{k}^{(q)}(\bm{x},\widetilde{\bm{x}})$ as
@@
Furthermore, for any $z\in[\eta_{1}2^{q-1},\min\{\eta_{1}2^{q},\eta_{2}\}]$, the following kernel function $\widehat{k}_{2}^{(q)}(\bm{x},\widetilde{\bm{x}};z)$ is positive definite (e.g., Lemma A.5 in [10]):
@@
By noting that the product of two positive definite kernels is also positive definite, the above expression implies that $2\mathbf{K}_{2}^{(q)}(\widetilde{\mathbf{X}}_{T},\widetilde{\mathbf{X}}_{T})-\widetilde{\mathbf{K}}_{2}^{(q)}(\mathbf{X}_{T},\mathbf{X}_{T})$ is the positive semi-definite matrix. Therefore, we have [^10] For any positive semi-definite matrices $A$, $B$ such that $A-B$ is positive semi-definite, we have $\lambda_{i}^{(A)}\geq\lambda_{i}^{(B)}$, where $(\lambda_{i}^{(A)})$ and $(\lambda_{i}^{(B)})$ is a non-negative eigenvalues of $A$ and $B$ with decreasing order. (This is a consequence of Courant–Fischer’s min-max theorem.) Therefore, we have $\det(A)=\prod_{i}\lambda_{i}^{(A)}\geq\prod_{i}\lambda_{i}^{(B)}=\det(B)$ for such $A$ and $B$.

### v3

@@ A.2 Proof of Lemma 5
By aggregating the conditions (70) and (77), we set $\overline{T}$ as the smallest natural number such that the following inequalities hold:
$\overline{T}\geq\left(\frac{e^{2}c_{d}\widetilde{C}_{\mathrm{SE}}}{2\ell^{2}c_{\mathrm{quad}}}\right)^{2}\beta_{T}\ln^{d+1}T,~~\mathrm{and}~~\overline{T}\geq\left(\frac{ec_{d}}{2c_{\mathrm{quad}}\ell^{2}}\right)^{4}\widetilde{C}_{\mathrm{SE}}^{4}\beta_{T}^{2}\ln^{2(d+1)}T.$ (80)
Then, from Eqs. (72) and (79), we have
$\sqrt{\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)}=O(\sqrt{\ln T}).$ (81)
@@ A.3 Proof of Lemma 6
From the upper bound of the discretization error in event $\mathcal{A}$, we have $\forall t\geq\sqrt{2/\varepsilon},\forall\bm{x}\in\mathcal{X},|f(\bm{x})-f([\bm{x}]_{t})|\leq\varepsilon/2$. Here, we set $\underline{\mathcal{T}}(\varepsilon)=\{t\in\mathbb{N}_{+}\mid t\geq\sqrt{2/\varepsilon}\}$. By relying on the standard argument of MIG [srinivas10gaussian], we observe the following inequality for any realizations and $\varepsilon>0$:
@@
where the second inequality follows from the definition of $\mathcal{T}(\varepsilon)$, and the last inequality follows from $\widetilde{t}\in\underline{\mathcal{T}}(\varepsilon)$ and event $\mathcal{A}$. Therefore, under $\mathcal{A}$, the inequality $\frac{\varepsilon}{2}+2\sqrt{\frac{C\beta_{\widetilde{t}}\gamma_{|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|}(\mathcal{X})}{|\mathcal{T}(\epsilon)\cap\underline{\mathcal{T}}(\varepsilon)|}}\geq 0$ must hold; otherwise, $\mu(\bm{x}_{\widetilde{t}};\mathbf{X}_{\widetilde{t}-1};\bm{y}_{\widetilde{t}-1})+\beta_{\widetilde{t}}^{1/2}\sigma(\bm{x}_{\widetilde{t}};\mathbf{X}_{\widetilde{t}-1})<\mu([\bm{x}^{\ast}]_{\widetilde{t}};\mathbf{X}_{\widetilde{t}-1};\bm{y}_{\widetilde{t}-1})+\beta_{\widetilde{t}}^{1/2}\sigma([\bm{x}^{\ast}]_{\widetilde{t}};\mathbf{X}_{\widetilde{t}-1})$, which contradicts $\bm{x}_{\widetilde{t}}\in\operatornamewithlimits{argmax}_{\bm{x}\in\mathcal{X}}\mu(\bm{x};\mathbf{X}_{\widetilde{t}-1};\bm{y}_{\widetilde{t}-1})+\beta_{\widetilde{t}}^{1/2}\sigma(\bm{x};\mathbf{X}_{\widetilde{t}-1})$. This further implies
@@ Appendix B Information Gain Upper Bound
Our analysis requires the upper bound of MIG with explicit dependence on the radius of the input domain. Several existing works [berkenkamp2019no, janz2022sequential, janz2020bandit] established such a result by extending the proof in [srinivas10gaussian]. However, the proof strategy in [srinivas10gaussian] result in $\widetilde{O}(T^{\frac{d(d+1)}{2\nu+d(d+1)}})$ upper bound of MIG in Matérn kernel, which is strictly worse than the best achievable $\widetilde{O}(T^{\frac{d}{2\nu+d}})$ upper bound. vakili2021information shows $\widetilde{O}(T^{\frac{d}{2\nu+d}})$ upper bound of MIG with $\nu>1/2$ under the uniform boundness assumption of the eigenfunctions. Furthermore, the following work [vakili2023kernelized] shows $\gamma_{T}(\{\bm{x}\in\mathbb{R}^{d}\mid\|\bm{x}\|_{2}\leq\eta\})=\widetilde{O}(\eta^{\frac{2\nu d}{2\nu+d}}T^{\frac{d}{2\nu+d}})$ for any radius $\eta>0$ if there exist eigenfunctions uniformly bounded without depending on $\eta>0$. Some of the related results supports the uniform boundness assumption under $d=1$ [janz2022sequential, yang2017frequentist], or under the approximated version of the original Matérn kernel [riutort2023practical, solin2020hilbert]; however, to our knowledge, we are not aware of any literature that rigorously support uniform boundness assumption under the general compact input domain with $d\geq 2$ and $\nu>1/2$. See Chapter 4.4 in [janz2022sequential] for the detailed discussion. Therefore, this section’s goal is twofold: (i) prove $\widetilde{O}(T^{\frac{d}{2\nu+d}})$ upper bound as claimed in [vakili2021information] without relying on the uniform boundness assumption, and (ii) clarify the explicit dependence on the input radius in the upper bound proved in (i).
@@
While the above results ignore the explicit dependence on $d$ and $\nu$, all the other parameters are explicitly stated in our upper bound of MIG. We would like to emphasize that we do not rely on the uniform boundness assumption of the eigenfunctions to prove the above results. In the above results for $k=k_{\mathrm{SE}}$, we obtain the same $O(\ln^{d+1}T)$ upper bound as that in [vakili2021information] except for the constant factor. For $k=k_{\mathrm{Mat\acute{e}rn}}$, we also obtain the same $\widetilde{O}(T^{\frac{d}{2\nu+d}})$ upper bound as that in [vakili2021information], while the logarithmic dependence get worse from $O\left(\ln^{d/(2\nu+d)}T\right)$ to $O\left(\ln^{(4\nu+d)/(2\nu+d)}T\right)$. Furthermore, the above result reveals the explicit dependence of the radius $\eta$ of the input domain. Regarding the case $k=k_{\mathrm{Mat\acute{e}rn}}$, our result suggests $\widetilde{O}(\eta^{\frac{2\nu d}{2\nu+d}}T^{\frac{d}{2\nu+d}})$ upper bound of MIG, which is consistent with that in [vakili2023kernelized] with uniform boundness assumption.
@@
The basic proof strategy follows that in [vakili2021information], which leverages the Mercer decomposition of the kernel. To bypass the uniform boundness assumption in the proof of [vakili2021information], we must resort to other specific properties of the eigenfunction. However, except for some exceptional cases, the eigenfunction of the kernel on the general compact domain is difficult to specify in an analytical form and complex to analyze. To avoid this issue, instead of studying the original definition of the MIG on $\mathbb{R}^{d}$, we consider reducing the original MIG on $\mathcal{X}\coloneqq\{\bm{x}\in\mathbb{R}^{d}\mid\|\bm{x}\|_{2}\leq 1\}$ to that on a hypersphere $\mathbb{S}^{d}\coloneqq\{\bm{x}\in\mathbb{R}^{d+1}\mid\|\bm{x}\|_{2}=1\}$ defined in $\mathbb{R}^{d+1}$. The eigensystems on $\mathbb{S}^{d}$ are one of the exceptional cases, whose eigenfunctions are specified as a special function on $\mathbb{S}^{d}$, called spherical harmonics [atkinson2012spherical, efthimiou2014spherical]. Indeed, by using the addition theorem of the spherical harmonics (Theorem 14), the existing works [iwazaki2024no, kassraie2022neural, vakili2021uniform] already demonstrated that the upper bound of MIG on $\mathbb{S}^{d}$ is proved as with [vakili2021information] without the uniform boundness assumption. We use their technique to show the upper bound of MIG under SE and Matérn kernels on $\mathbb{R}^{d}$, while the original motivation of these existing works is to study the MIG under the neural tangent kernel on a hypersphere. The remaining parts of this section are constructed as follows:
@@
  • In Section B.3, we provide the general upper bound of the information gain on $\mathbb{S}^{d}$ (Lemma 15), represented by the kernel function’s eigenvalues. This subsection’s result has no intrinsic change from those in [kassraie2022neural, vakili2021uniform]; however, we provide details for completeness.
@@ B.1 Reduction of the MIG on $\mathbb{R}^{d}$ to $\mathbb{S}^{d}$
where $\widetilde{\mathbf{K}}]_{i,j}=k(\widetilde{\bm{x}}_{i},\widetilde{\bm{x}}_{j})/k(\bm{x}_{i},\bm{x}_{j})$. Here, $A\odot B$ denotes the Hadamard product of the matrices $A$ and $B$. Then, Oppenheim inequality (e.g., Theorem 7.27 in [zhang2011matrix]) implies $\det(A\odot B)\geq\det(B)\prod_{i}A_{ii}$ for any positive semi-definite matrices $A$ and $B$. Therefore, if $\widetilde{\mathbf{K}}$ is a positive semi-definite matrix, Eq. (114) immediately implies
@@
Similarly to the proof for $k=k_{\mathrm{SE}}$, we consider the application of Oppenheim inequality; however, the positive semi-definiteness of element-wise quotient matrix $\widetilde{\mathbf{K}}$ is unknown for $k=k_{\mathrm{Mat\acute{e}rn}}$. To avoid this problem, we leverage the following representation of $k_{\mathrm{Mat\acute{e}rn}}$, which is given as the form of the lengthscale mixture of the SE kernel [tronarp2018mixture]:
@@
[…]
with some $\eta_{2}>\eta_{1}>0$ [^8] Note that the linear combination of the positive definite kernel with non-negative coefficients and its limit are also positive definite. Therefore, $k_{1},k_{2}$, and $k_{3}$ are also positive definite as far as $\eta_{2}>\eta_{1}$.. Then, as with the proof of Theorem 3 in [krause2011contextual], we have
@@
[…]
@@
[…]
@@
[…]
@@
Furthermore, for any $z\in[\eta_{1}2^{q-1},\min\{\eta_{1}2^{q},\eta_{2}\}]$, the following kernel function $\widehat{k}_{2}^{(q)}(\bm{x},\widetilde{\bm{x}};z)$ is positive definite (e.g., Lemma A.5 in [capone2022gaussian]):
@@
By noting that the product of two positive definite kernels is also positive definite, the above expression implies that $2\mathbf{K}_{2}^{(q)}(\widetilde{\mathbf{X}}_{T},\widetilde{\mathbf{X}}_{T})-\widetilde{\mathbf{K}}_{2}^{(q)}(\mathbf{X}_{T},\mathbf{X}_{T})$ is the positive semi-definite matrix. Therefore, we have [^10] For any positive semi-definite matrices $A$, $B$ such that ABA-B is positive semi-definite, we have $\lambda_{i}^{(A)}\geq\lambda_{i}^{(B)}$, where $(\lambda_{i}^{(A)})$ and $(\lambda_{i}^{(B)})$ is a non-negative eigenvalues of $A$ and $B$ with decreasing order. (This is a consequence of Courant–Fischer’s min-max theorem.) Therefore, we have $\det(A)=\prod_{i}\lambda_{i}^{(A)}\geq\prod_{i}\lambda_{i}^{(B)}=\det(B)$ for such $A$ and $B$.